# LLaVA-3D: A Simple yet Effective Pathway to Empowering LMMs with 3D-awareness

### v1

@@ Abstract
Recent advancements in Large Multimodal Models (LMMs) have greatly enhanced their proficiency in 2D visual understanding tasks, enabling them to effectively process and understand images and videos. However, the development of LMMs with 3D-awareness for 3D scene understanding has been hindered by the lack of large-scale 3D vision-language datasets and powerful 3D encoders. In this paper, we introduce a simple yet effective framework called LLaVA-3D. Leveraging the strong 2D understanding priors from LLaVA, our LLaVA-3D efficiently adapts LLaVA for 3D scene understanding without compromising 2D understanding capabilities. To achieve this, we employ a simple yet effective representation, 3D Patch, which connects 2D CLIP patch features with their corresponding positions in 3D space. By integrating the 3D Patches into 2D LMMs and employing joint 2D and 3D vision-language instruction tuning, we establish a unified architecture for both 2D image understanding and 3D scene understanding. Experimental results show that LLaVA-3D converges 3.5$\times$ faster than existing 3D LMMs when trained on 3D vision-language datasets. Moreover, LLaVA-3D not only achieves state-of-the-art performance across various 3D tasks but also maintains comparable 2D image understanding and vision-language conversation capabilities with LLaVA.
Figure 1: Overview of LLaVA-3D. Block (a) shows that LLaVA-3D could perform both 2D and 3D vision-language tasks. The left block (b) shows that compared with previous 3D LMMs, our LLaVA-3D achieves state-of-the-art performance across a wide range of 3D benchmarks while maintaining a comparable performance on various 2D benchmarks compared with LLaVA-1.5. The middle block (c) demonstrates that LLaVA-3D is built on the 2D LMM: LLaVA, and leverages 3D patches to endow it with 3D spatial awareness, enabling it to perform various 3D vision-and-language tasks in the physical world. The right blocks (d) and (e) highlights the significantly faster convergence and inference speeds of LLaVA-3D compared to existing 3D LMMs.
@@ 1 Introduction
Recent advancements in Large Multimodal Models (LMMs) [2, 8, 30, 43] have significantly enhanced their ability to understand and reason over visual and language inputs, leading to remarkable performance in 2D visual tasks, such as 2D perception, understanding, and reasoning. Despite their advanced perceptual and reasoning capabilities, LMMs are primarily confined to virtual interactions through images or video, lacking the critical ability to interact with the physical world. To enable their deployment in real-world applications and to facilitate the emergence of new capabilities through physical interactions, it is imperative to equip LMMs with 3D spatial intelligence.
@@
A key aspect of 3D spatial intelligence is the ability to perceive and understand the 3D world. Similar to how 2D LMMs align 2D visual features with language models using large-scale 2D vision-language datasets, a common approach to developing 3D LMMs [12, 20, 19] involves integrating 3D features encoded from point clouds into Large Language Models (LLMs) and training them on 3D point cloud-language datasets. However, in contrast to the abundance of large-scale 2D datasets, 3D datasets remain relatively scarce. Meanwhile, there are no powerful pre-trained 3D point cloud encoders, akin to CLIP ViT [42] in 2D, to serve as a foundational model that can provide strong 3D features to LLMs.
@@
Since real-world embodied agents typically rely on ego-centric, multi-view images as their raw observations, we aim to build a 3D foundation model based on such inputs rather than 3D point clouds. There have been attempts [18, 17] to leverage the 2D foundation models, like CLIP, alongside LLMs to advance this goal. These methods resort to 2D object segmentation results [26] to extract and aggregate CLIP features from object-centric image patches, constructing pixel-aligned 3D scene features [23]. However, this multi-view image object segmentation and object-centric feature extraction pipeline is inherently complex and computationally intensive. In contrast, 2D LMMs [8, 30, 37, 38, 2] directly leverage CLIP’s image patch features with richer, fine-grained information for effective image understanding and visual reasoning. This naturally leads to the question: Can we directly build a 3D LMM upon the strong 2D priors from 2D LMMs, bypassing the obstacles in 3D data scale and 3D encoders?
@@
In light of recent progress in 2D LMMs, we propose a simple yet effective framework, LLaVA-3D, which extends the well-established 2D LLaVA model to efficiently comprehend the 3D world while preserving its robust 2D multimodal perception and reasoning capabilities. The core innovation in our approach is the introduction of 3D Patch, a new 3D representation that bridges 2D features within a 3D spatial context. This representation is derived by simply augmenting 2D patch-wise features with 3D positional embeddings, eliminating the need for additional complex processing. After enhancing the 2D visual tokens to 3D patch features, we explore various pooling strategies to compress the 3D patch features across extensive input frames before being fed into LLM. Fine-tuned on the existing 3D vision-language datasets, our model converges rapidly and acquires 3D spatial understanding and grounding capabilities. Furthermore, the unified model architecture allows LLaVA-3D to retain the strong 2D understanding and reasoning abilities of LLaVA through joint instruction-tuning on 2D vision-language datasets.
Our experimental results demonstrate that LLaVA-3D achieves state-of-the-art performance on a wide range of 3D tasks and benchmarks [3, 40, 39, 41, 13, 51, 6], including 3D captioning, 3D question answering, and 3D grounding while requiring significantly less training time and fewer epochs than existing 3D LMMs. Additionally, LLaVA-3D achieves comparable capabilities in 2D image understanding, reasoning, and conversation to LLaVA through joint tuning on 2D and 3D vision-language instructions.
@@ 2 Related Work
Building on the success of recent LLMs, numerous studies [37, 38, 2, 30, 8, 35] explored LMMs that can jointly process visual and linguistic information. For example, LLaVA [37, 38] aligned 2D images with language models through an image encoder and a projection layer, while BLIP2 [30] employed a more sophisticated Q-Former architecture to guide the compression of visual features using textual cues. However, most early 2D LMMs were trained on single-image datasets, limiting their ability to tackle the complexities of multi-image understanding. Recently, there has been increasing interest in expanding LMMs to handle multi-image inputs, addressing the demands of real-world scenarios. For video LMMs [34, 31, 47, 28], multi-image input forms the basis for capturing temporal or action-related dynamics across sequences of video frames. On the other hand, multi-view images of the 3D scene can implicitly reveal 3D spatial relationships and other abstract relations in the environment. Recent works  [41, 36] explored whether 2D LMMs [1, 44] can leverage multi-view images to perform spatial understanding. However, these methods primarily relied on implicit learning from the data, without directly modeling the 3D world. In contrast, our LLaVA-3D explicitly models the 3D world from multi-view images, enabling advanced 3D spatial understanding and grounding capabilities.
Injecting 3D into LLMs.
@@
As 2D LMMs achieved substantial progress in visual perception, similar efforts have been made in the 3D domain. For 3D scene-level understanding, recent works explored ways to integrate 3D inputs such as point clouds [19, 12, 20] or multi-view images [18, 17] into LLMs to enable advanced 3D scene understanding and reasoning. An important distinction among these methods is how they construct the 3D scene representation. LL3DA [12] directly used a scene-level 3D point cloud encoder to extract the 3D scene representation. LEO [20] and Chat3D v2 [19] first segmented 3D objects from the scene point cloud using the off-the-shelf 3D instance segmentation model, and then independently extracted 3D object features with object-level 3D encoders to represent the 3D scene. On the other hand, starting from multi-view images, 3D-LLM [18] and Scene-LLM [17] resorted to manually crafted 2D object segmentation to extract and aggregate CLIP features from object-centric image patches, constructing pixel-aligned 3D point representation. Unlike these approaches, our LLaVA-3D directly builds on the well-trained 2D LMM with multi-view images as input. Utilizing the 3D position embeddings, it brings the the 2D patches within a 3D spatial context to construct 3D Patches. This 3D representation enables quick adaption of LLaVA for 3D scene understanding while preserving its strong 2D image understanding ability.
@@ 3 Method
Figure 2: LLaVA-3D Architecture. Based on LLaVA, we directly add the corresponding 3D position embeddings to 2D patch visual tokens of multi-view images to construct the 3D Patches, then the 3D Patches will undergo 3D pooling and be sent into the projection layer of LLaVA to map into the LLM space and align with the LLM using 3D-visual-language data.
@@ 3.1 Preliminary
We choose LLaVA [38], a 2D LMM, to explore building a 3D LMM based on it. LLaVA uses the pre-trained CLIP encoder to extract the 2D patch features $X_{v}\in\mathbb{R}^{c\times w\times h}$ from the input image $X\in\mathbb{R}^{3\times W\times H}$, and then align the 2D patch features $X_{v}$ into with LLM space with the projection layer. A simple multi-view image adaptation [29] for LLaVA could be extracting multi-view 2D Patch features from multi-view images $X^{\prime}_{v}\in\mathbb{R}^{V\times c\times w\times h}$ and sequentially feeding them into LLM for multi-image understanding. To empower LLaVA with 3D awareness, we introduce 3D Patch, a novel 3D representation that integrates 3D spatial information into 2D patch features.
@@
Our 3D Patch representations are built upon the 2D patch features $X^{\prime}_{v}$ extracted from multi-view images with CLIP visual encoder to leverage the strong visual-semantic alignment. To construct the 3D Patches, we inject the 3D position information into the aforementioned 2d patches so that the 3D Patches can explicitly model 3D spatial information while preserving the semantic information from 2D patches. As illustrated in left block of Fig. 2, given the multi-view 2D patch features $X^{\prime}_{v}\in\mathbb{R}^{V\times c\times w\times h}$, we obtain their 3D positions $P\in\mathbb{R}^{V\times 3\times w\times h}$ in the 3D world, using nearest neighbor depth and known camera intrinsic and extrinsic parameters. The 3D positions $P$ are then encoded into 3D position embeddings $P^{\prime}\in\mathbb{R}^{V\times w\times h\times d}$ through a learnable two-layer MLP, which are subsequently added to the 2D patch visual tokens, resulting in the 3D patches $X^{\prime}_{3D}\in\mathbb{R}^{V\times w\times h\times d}$:
$X^{\prime}_{3D}=X^{\prime}_{v}+\text{MLP}(P^{\prime})$ (1)
@@
While the 3D Patches equip 2D patches with 3D spatial awareness, the number of 3D patches scales directly with the number of input images. Capturing a full 3D scene often necessitates a large set of images, which significantly increases the computational overhead for large language models (LLMs). To address this, we introduce a 3D-aware pooling mechanism to reduce the number of 3D patches, as illustrated in the middle block of Fig. 2.
In the 2D image or video domain, pooling is commonly applied along the 2D spatial or temporal dimensions to compress the number of tokens and extract essential semantic information. However, for 3D scene understanding, we must pool the 3D patches based on their 3D locations to ensure these features can cover and preserve the entire scene’s structure as completely as possible. We explore two parameter-free pooling strategies to achieve this:
Voxelization Pooling.
Voxelization discretizes the 3D space into a volumetric grid. Within each occupied voxel, the 3D patches undergo average pooling, resulting in updated voxel visual tokens. Only the visual tokens from occupied voxels are passed to the LLM, with the number of tokens varying across different 3D scenes. While the number of 3D patches scales with the number of images, the number of voxel patches is only determined by the partition of voxel grids. We can easily balance the number of visual tokens and the preservation of fine-grained scene features by controlling the voxel size.
FPS Pooling.
Farthest Point Sampling (FPS) is a widely used sampling strategy to select a representative subset of points from a larger set of points cloud. We apply FPS to sample 3D patches from multi-view images to a fixed number of tokens, ensuring that the sampled tokens represent the entire scene structure. While fixing the number of tokens helps the LLM efficiently process visual information, it may also result in loss of scene information.
Compared to FPS Pooling, Voxelization Pooling offers equivalent efficiency in visual token compression while preserving more detailed scene information. Furthermore, the explicit construction of voxel grids can better handle dynamic 3D scene updates, whereas FPS Pooling excels at preserving the overall 3D scene structure. We conduct quantitative experiments in Sec. 5.7 to thoroughly assess the effectiveness of these pooling strategies.
@@
In the previous sections, we detailed the construction of the 3D scene representation from multi-view images, establishing the foundation for further interaction with the 3D scene. Building on this, the LLM could process multi-modal inputs such as the 3D scene, language instructions, and 3D coordinate cues to generate outputs such as language responses and 3D bounding boxes, as illustrated in the right block of Fig. 2. In this section, we introduce how the model is equipped to interpret 3D coordinate information from inputs and subsequently output precise 3D bounding boxes when specific location-related task requirements are needed.
Encoding of 3D Coordinate Input.
In scenarios such as 3D dense object captioning or object-centric question answering, the language instruction contains 3D coordinates. To handle such tasks, we introduce the 3D Coordinate Token to allow the model to integrate the provided coordinates as context into its reasoning processes. Specifically, we obtain the 3D coordinate token by feeding the 3D coordinates through a two-layer MLP. These 3D coordinate tokens are fed into LLM together with 3D Patch tokens and text tokens, enabling 3D-aware perception and reasoning.
Decoding of 3D Bounding Box Output.
The integration of the 3D coordinate token empowers the model to process 3D coordinate information from input instructions effectively. However, experiments show that directly outputting 3D bounding boxes is quite challenging for LLM, and empirically, the performance is poor. To overcome this, we leverage an approach similar to previous method  [51], introducing a specialized location token that guides the grounding module to generate accurate 3D bounding boxes. Specifically, the LLM predicts a special location token to represent a 3D box prediction when the task necessitates 3D bounding box outputs. We then derive the last layer embedding of this location token and send it into the grounding module. The grounding module utilizes the location token embedding and the 3D scene feature to predict the 3D box coordinates of the target object. This process facilitates the precise generation of 3D bounding box outputs. More details can be found in the appendix.

### v3

# LLaVA-3D: A Simple yet Effective Pathway to Empowering LMMs with 3D Capabilities

@@ Abstract
Recent advancements in Large Multimodal Models (LMMs) have greatly enhanced their proficiency in 2D visual understanding tasks, enabling them to effectively process and understand images and videos. However, the development of LMMs with 3D scene understanding capabilities has been hindered by the lack of large-scale 3D vision-language datasets and powerful 3D encoders. In this paper, we introduce a simple yet effective framework called LLaVA-3D. Leveraging the strong 2D visual understanding priors from LLaVA, our LLaVA-3D efficiently adapts LLaVA for 3D scene understanding without compromising 2D understanding capabilities. To achieve this, we utilize the 3D position embeddings to enhance the 2D CLIP Patches with 3D spatial context information and construct 3D patches. By integrating the 3D position embeddings into 2D LMMs and employing joint 2D and 3D vision-language instruction tuning, we establish a unified architecture for both 2D visual understanding and 3D scene understanding. In contrast to previous 3D LMMs, LLaVA-3D supports decoding accurate 3D spatial perception outputs, e.g., 3D bounding boxes, directly from these 3D patches, without relying on the time-consuming off-the-shelf 3D segmentors. Experimental results show that LLaVA-3D converges 3.5$\times$ faster than existing 3D LMMs when trained on 3D vision-language datasets. Moreover, LLaVA-3D not only achieves state-of-the-art performance across various 3D tasks but also maintains comparable 2D visual understanding and vision-language conversation capabilities with LLaVA.
Figure 1: Overview of LLaVA-3D. The left block (b) shows that our LLaVA-3D achieves state-of-the-art performance across various 3D scene understanding benchmarks. Notably, LLaVA-3D maintains the comparable performance on 2D multimodal benchmarks compared to LLaVA-Video [53]. The middle block (c) demonstrates that LLaVA-3D leverages 3D patches to endow the 2D LMMs with 3D spatial awareness, enabling it to perform various 3D vision-and-language tasks in the physical world. The right block (d) and (e) highlight the significantly faster convergence and inference speeds of LLaVA-3D compared to existing 3D LMMs.
@@ 1 Introduction
Recent advancements in Large Multimodal Models (LMMs) [2, 5, 27, 45] have significantly enhanced their ability to understand and reason over visual and language inputs, leading to remarkable performance in 2D visual tasks. Despite their advanced perceptual and reasoning capabilities, LMMs are primarily confined to virtual interactions through images or video, lacking the critical ability to interact with the physical world. To enable their deployment in real-world applications and to facilitate the emergence of new capabilities through physical interactions, it is imperative to equip LMMs with 3D spatial intelligence.
@@
A key aspect of 3D spatial intelligence is the ability to perceive and understand the 3D world. Similar to how 2D LMMs align 2D visual features with language models using large-scale 2D vision-language datasets, a common approach to developing 3D LMMs [8, 18, 16] involves integrating 3D features encoded from point clouds into Large Language Models (LLMs) and training them on 3D point cloud-language datasets. However, in contrast to the abundance of large-scale 2D datasets, 3D datasets remain relatively scarce. Meanwhile, there are no powerful pre-trained 3D point cloud encoders, akin to CLIP ViT [43] in 2D, to provide strong and generalizable 3D features to LLMs.
@@
Since real-world embodied agents typically rely on ego-centric, multi-view images as raw observations, we aim to build a 3D foundation model based on such inputs rather than 3D point clouds. There have been attempts [15, 14] to leverage the 2D foundation models, like CLIP, alongside LLMs to advance this goal. These methods resort to 2D object segmentation results [24] to extract and aggregate CLIP features from object-centric image patches, constructing pixel-aligned 3D scene features [22]. However, this pipeline is inherently complex and computationally intensive. In contrast, 2D LMMs [5, 27, 33, 34, 2, 26, 53] directly leverage CLIP’s image patch features with richer, fine-grained information for effective 2D understanding and reasoning. This naturally leads to the question: Can we directly build a 3D LMM upon the strong 2D priors from 2D LMMs, bypassing the obstacles in 3D data scale and 3D encoders?
@@
In light of recent progress in 2D LMMs, we propose a simple yet effective framework, LLaVA-3D, which extends the well-established LLaVA model to efficiently comprehend the 3D world while preserving its robust 2D multimodal perception and reasoning capabilities. Inspired by ODIN [21], which leverages the positional encodings for unified 2D and 3D segmentation, our LLaVA-3D encodes the 3D spatial coordinates into 3D position embeddings, and incorporate them into the 2D CLIP patches in LLaVA to construct 3D patches. These 3D patches then undergo adaptive token compression via efficient pooling strategies before LLM processing. To further adapt LLaVA for tasks involving accurate 3D information in input or output, we devise an effective 3D-aware position encoding and decoding approach based on the 3D patches, without relying on the off-the-shelf 3D segmentors used in previous works [16, 17, 18]. Fine-tuned on the existing 3D vision-language datasets, our model converges rapidly and acquires 3D spatial understanding and grounding capabilities. Furthermore, the unified model architecture allows LLaVA-3D to retain the strong 2D understanding and reasoning abilities of LLaVA through joint instruction-tuning on 2D vision-language datasets.
Although LLaVA-3D is first built upon the LLaVA-family LMMs, it is a general extension method with simple yet effective designs for equipping any 2D LMM with 3D capabilities. With preliminary attempts based on different 2D LMMs [34, 53, 12] as the foundation model, we empirically observe that LLaVA-3D can benefit from stronger perception and understanding capabilities from large-scale 2D multi-modal pretraining, especially the video-based pretraining given the intrinsic connection between video and multi-view 3D scene representations. As illustrated in Fig. 1, experimental results demonstrate that LLaVA-3D achieves state-of-the-art performance on a wide range of 3D scene understanding benchmarks [3, 39, 38, 40, 10, 56], covering tasks such as 3D dense captioning and 3D question answering. Notably, thanks to the minimal 3D designs and powerful pretraining, LLaVA-3D attains these results with significantly less training time and fewer epochs than existing 3D LMMs, without large-scale data and pretraining costs for alignment. It could also achieve promising 3D visual grounding results without relying on the time-consuming offline 3D object preprocessing [16, 17, 18]. Furthermore, LLaVA-3D maintains capabilities comparable to state-of-the-art 2D LMMs in 2D visual understanding, reasoning, and conversation through joint tuning on 2D and 3D vision-language instructions.
@@ 2 Related Work
2D LMMs. Building on the success of recent LLMs, numerous studies [33, 34, 2, 27, 5, 31] explored LMMs that can jointly process visual and linguistic information. For example, LLaVA [33, 34] aligned 2D images with language models through an image encoder and a projection layer, while BLIP2 [27] employed a sophisticated Q-Former architecture to guide the compression of visual features using textual cues. However, most early 2D LMMs were trained on single-image datasets, limiting their ability to tackle multi-image understanding. Recently, there has been increasing interest in expanding LMMs to handle multi-image inputs, addressing the demands of real-world scenarios. For video LMMs [30, 28, 50, 26], multi-image input forms the basis for capturing temporal or action-related dynamics across sequences of video frames. On the other hand, multi-view images of the 3D scene can implicitly reveal 3D spatial relationships and other abstract relations in the environment. Recent works  [40, 32] explored whether 2D LMMs [1, 46] can leverage multi-view images to perform spatial understanding. However, these methods primarily relied on implicit learning from the data, without directly modeling the 3D world. In contrast, our LLaVA-3D explicitly models the 3D world from multi-view images, enabling advanced 3D spatial understanding and grounding capabilities.
@@
Injecting 3D into LLMs. As 2D LMMs achieved substantial progress in visual perception, similar efforts have been made in the 3D domain. For 3D scene-level understanding, recent works explored ways to integrate 3D inputs such as point clouds [16, 8, 18, 17] or multi-view images [15, 14, 55, 42] into LLMs to enable advanced 3D scene understanding and reasoning. An important distinction among these methods is how they construct the 3D scene representation. LL3DA [8] directly used a scene-level 3D point cloud encoder to extract the 3D scene representation. LEO [18] and Chat-Scene [17] first segmented 3D objects from the scene point cloud using the off-the-shelf 3D instance segmentation model and then independently extracted 3D object features with object-level 3D encoders to represent the 3D scene. On the other hand, starting from multi-view images, 3D-LLM [15] and Scene-LLM [14] resorted to manually crafted 2D object segmentation to extract and aggregate CLIP features from object-centric image patches, constructing pixel-aligned 3D point representation. Unlike these approaches, our LLaVA-3D directly builds on the well-trained 2D LMM with multi-view images as input. Utilizing the 3D position embeddings, it brings the 2D patches within a 3D spatial context to construct 3D Patches. This 3D scene representation enables quick adaption of LLaVA for 3D scene understanding while preserving its strong 2D image understanding ability.
Joint Modeling of 2D and 3D. Recent works [35, 36, 21] explored leveraging existing 2D foundation models to enhance 3D perception for detection and segmentation tasks. These methods extract 2D features from multi-view images using 2D foundation models, and then construct 3D position-aware features by incorporating 3D position embeddings for improved 3D detection and segmentation. ODIN [21] utilized the posed RGB-D images as input for 3D instance segmentation. It leverages the powerful 2D pre-trained backbone and differentiates between 2D and 3D features by using distinct learnable position encodings, with 2D features represented by pixel coordinates and 3D features represented by 3D coordinates. This unified architecture facilitated joint training on both 2D and 3D datasets, further enhancing the performance of 3D segmentation. Our LLaVA-3D first integrates the 3D position-aware features into 2D LMMs, enabling 2D LMMs to achieve 3D understanding. Analogous to ODIN [21], this modeling approach enables joint 2D-3D training, allowing the model to process and reason about both 2D and 3D tasks in a unified framework.
@@ 3 Method
Figure 2: LLaVA-3D Architecture. Based on LLaVA-Video, we directly add the corresponding 3D position embeddings to 2D patch visual tokens of multi-view images to construct the 3D patches. Considering the context length support of the base model, we skip the 3D Patch Pooling stage and directly send the 3D patches into the 3D-aware Position Encoding and Decoding process to perform various 3D understanding tasks.
@@ 3.1 Preliminary
We choose LLaVA-Video [53] as the base model to build the 3D LMM. For each frame image, LLaVA-Video uses the pre-trained CLIP encoder to split the image $X\in\mathbb{R}^{3\times W\times H}$ into patches at the patch size $P$ and extract the 2D patch features $X^{\prime}\in\mathbb{R}^{c\times w\times h}$, where $h=\left\lfloor\frac{H}{P}\right\rfloor,w=\left\lfloor\frac{W}{P}\right\rfloor$, and then align the 2D patch features $X_{v}$ into with LLM space with the projection layer. For the 3D scene understanding, the multi-view image patch features $X^{\prime}_{v}\in\mathbb{R}^{V\times c\times w\times h}$ are sequentially sent into LLM. To empower LLaVA-Video with 3D capabilities, we incorporate the 3D position embeddings into 2D patches to obtain the 3D patches.
@@
Our 3D patch representations are built upon the 2D patch features $X^{\prime}_{v}$ extracted from multi-view images with CLIP visual encoder to leverage the strong visual-semantic alignment. To construct the 3D patches, we inject the 3D position information into the 2D patches so that the 3D patches can explicitly model 3D spatial information while preserving the semantic information from 2D patches. As illustrated in left block of Fig. 2, given the multi-view 2D patch features $X^{\prime}_{p}\in\mathbb{R}^{V\times d\times w\times h}$ after the projection layer, we obtain their 3D positions $P\in\mathbb{R}^{V\times 3\times w\times h}$ in the 3D world, using nearest neighbor depth and known camera intrinsic and extrinsic parameters, following ODIN [21]. The 3D positions $P$ are then encoded into 3D position embeddings $P^{\prime}\in\mathbb{R}^{V\times d\times w\times h}$ through the 3D Position Encoding Layer which consists of a learnable two-layer MLP. The 3D position embeddings are subsequently added to the 2D patch visual tokens, resulting in the 3D patches $X^{\prime}_{3D}\in\mathbb{R}^{V\times d\times w\times h}$:
$X^{\prime}_{3D}=X^{\prime}_{p}+P^{\prime}$ (1)
@@
While 3D patches enhance 2D patches with spatial information, they increase linearly with the number of input images and may exceed the context length of LLM. To address this, we introduce a 3D-aware pooling mechanism to reduce the number of 3D patches when token compression is needed, as illustrated in the middle block of Fig. 2.
In the 2D image or video domain, pooling is commonly applied along the 2D spatial or temporal dimensions to compress the number of tokens and extract essential semantic information. However, for 3D scene understanding, we pool the 3D patches based on their 3D locations to ensure these features can cover and preserve the entire scene’s structure as completely as possible. We explore two parameter-free pooling strategies to achieve this:
Voxelization Pooling. Voxelization discretizes the 3D space into a volumetric grid, with 3D patches undergoing average pooling within each occupied voxel, resulting in updated voxel visual tokens. Only the visual tokens from the occupied voxels are passed to the LLM, and the number of tokens varies across different 3D scenes. While the number of 3D patches scales with the number of images, the number of voxel tokens depends solely on the partitioning of the voxel grid. By adjusting the voxel size, we can effectively balance the trade-off between the number of visual tokens and the preservation of fine-grained scene features.
FPS Pooling. Farthest Point Sampling (FPS) is a widely used sampling strategy [20, 37] to select a representative subset of points from a larger set of points cloud. We apply FPS to sample 3D patches from multi-view images to a fixed number of tokens, ensuring that the sampled tokens represent the entire scene structure. While fixing the number of tokens helps the LLM efficiently process visual information, it may also result in loss of scene information.
@@
In the previous sections, we detailed the construction of the 3D scene representation from multi-view images, establishing the foundation for further interaction with the 3D scene. Building on this, the LLM could process multi-modal inputs such as the 3D scene, language instructions, and 3D coordinate cues to generate outputs such as language responses and 3D bounding boxes, as illustrated in the right block of Fig. 2. In this section, we introduce how the model is equipped to interpret accurate 3D coordinate information from inputs and subsequently output precise 3D bounding boxes when specific location-related task requirements are needed.
Encoding of 3D Coordinate Input. In scenarios such as 3D dense object captioning or object-centric question answering, the language instruction contains 3D coordinates. To handle such tasks, we introduce the 3D Coordinate Token to allow the model to integrate the provided coordinates as context into its reasoning processes. Specifically, we obtain the 3D coordinate token by feeding the 3D coordinates through the 3D position encoding layer. The 3D coordinate tokens are fed into LLM together with 3D patch tokens and text tokens, enabling 3D coordinate-aware perception and reasoning.
Decoding of 3D Bounding Box Output. The integration of the 3D coordinate token enables the model to process 3D coordinate information from instructions effectively. However, experiments reveal that directly outputting 3D object location coordinates is non-trivial for the LLM, resulting in notably poor performance in 3D visual grounding. To handle this, previous 3D LMMs [17, 55, 16] tend to rely on the offline time-consuming 3D object extraction method and convert the task to a 3D object selection task, which is not suitable for real world application. In the section, we introduce the efficient Grounding Decoder: The process begins with a set of instance queries sampled via Farthest Point Sampling from the 3D patches. The 3D patches contain rich language-aligned semantic information but lack detailed geometric structure. To handle the grounding task, the grounding decoder guides instance queries to capture the geometry awareness from the 3D patches and aggregate information from the LLM. Specifically, in each decoder layer, we conduct cross-attention between the instance queries and the 3D patch features and then concatenate the updated queries and location token for distance-adaptive self-attention [13] to capture the relative relationship. To model the local object geometry information at different scales and reduce the computational complexity, we apply multi-scale 3D k-NN attention with relative 3D positional embeddings when cross-attending to 3D patch features. The updated instance queries are sent to the grounding head to predict corresponding 3D bounding boxes, and the matching score is calculated based on the similarity of the queries and location token. More details are provided in our supplementary material.